IF A BOY WALKS AT A SPEED OF 4 KM/HR,HE REACHES HIS SCHOOL 10 MINUTES BEFORE THE SCHOOL TIME.IF HE WALKS 3 KM/HR HE REACHES THE SCHOOL 10 MINUTES LATE.FIND THE DISTANCE BETWEEN HIS HOUSE AND HIS SCHOOL.
Answer by stanbon(75887)

IF A BOY WALKS AT A SPEED OF 4 KM/HR,HE REACHES HIS SCHOOL 10 MINUTES BEFORE THE SCHOOL TIME.IF HE WALKS 3 KM/HR HE REACHES THE SCHOOL 10 MINUTES LATE.FIND THE DISTANCE BETWEEN HIS HOUSE AND HIS SCHOOL.
----------------------
To-school DATA:
distance = x km ; rate = 4 kph ; time = x/4 hrs
----
From school DATA:
distance = x km ; rate = 3 kph ; time = x/3 hrs
------------------
Note the time difference is 25 minutes.
------------------
Equation:
from time - to time = 25/60 hrs
---
x/3 - x/4 = 25/60
---
Multiply thru by 60 to get:
20x - 15x = 25
5x = 25
x = 5 km (distance from house to school)
=========================================
Cheers,
Stan H.